OpenStudy (anonymous):

how do i find sin(3t) for all degrees on the unit circle?

OpenStudy (anonymous):

Plug in all the degrees on the unit circle and see what you get for each one.

OpenStudy (anonymous):

okay, i did that

OpenStudy (anonymous):

Really? All 360 of them? That must have taken you a while.

OpenStudy (anonymous):

I'm sorry, not all. I have 30 and 45 degree angles

OpenStudy (anonymous):

if you find them all in the range 0 to 90, then you can find the sine of the other angles by simple symmetry

OpenStudy (anonymous):

Okay. Find sin(3(30)) and sin(3(45)).

OpenStudy (anonymous):

sin(45) is easy...sin(30) is easy...so is sin(60), sin(90) and sin(0) sin(15) = sin(45 - 30) = sin45cos30 - cos45sin30 = sqrt(2)/4(sqrt(3) - 1) sin(75) = sin(45 + 30) = sin45cos30 + cos45sin30 = sqrt(2)/4(sqrt(3) + 1)

OpenStudy (anonymous):

Okay, here is what i did. I am making a graph to show f(t)= sin(t) + sin(3t). I have 5 columns; first has the angles 30, 45 degrees, second column has (t) in radians, third column has sin(t) in decimals. Now I am trying to fill in my fourth column with sin (3t). then my fifth column will have sin(t) + sin (3t)

OpenStudy (anonymous):

second column: pi/6, pi/4 third column: 0.5, 0.7071 fourth column: sin(90) = 1, sin(135)=sin(45)=0.7071 fifth column: 0.5 + 1 = 1.5, 0.7071 + 0.7071 = 1.4142

OpenStudy (anonymous):

fourth needs sin (3t) would that be sin times 3

OpenStudy (anonymous):

sin(3t) is the sine of 3 times the angle t so if t = 30, sin(3t) = sin(90)

OpenStudy (anonymous):

Oh, okay. Thank you

OpenStudy (anonymous):

Just to make sure I am correct sin(3t) of 30=1, 45=.7, 60=-1 90=.7, 120=0